References for Peter Sarnak’s course at AWS 2001

SURVEY:

1. Iwaniec and Sarnack, “Perspectives on the analytic theory of Lfunctions,” to appear in GAFA (2001).

SUBCONVEXITY:

2. Friedlander and Iwaniec, “A mean value theorem for character sums,” Michigan Math Jnl 39, 153-159 (1992).

3. Duke, Friedlander and Iwaniec, “Bounds for automorphic L-functions I and II,” Invent. Math 112, 1-8 (1993); Invent. Math 115, 219-239 and 209-217 (1994).

4. Sarnak, “Estimates for Rankin Selberg L-functions and quantum ergodicity,” preprint 2001.

5. Sarnak, “Integrals of products of eigenfunctions,” Imrn 6, 251-260 (1994).

SPECIAL VALUES OF L-FUNCTIONS:

6. Waldspurger, “Sur les coefficients de Fourier – demi enitier,” J. Math. Pures et Appl. 60, 365-384 (1981).

7. T. Watson, “Rankin triple products and quantum chaos,” Thesis Princeton U (2001).

NON-VANISHING AND MOLLIFICATION:

8. Iwaniec and Sarnak, “The non-vanishing of central values of automorphic L-functions and Landau-Siegel zeros,” Israel Math Jnl, Feb. 2001.

9. Kowalski, Michel and Vanderkam, “Mollification of the fourth moments of automorphic L-functions and arithmetic applications,” Invent. 142, 95-151 (2000).

PETERSSON FORMULA:

10. Iwaniec, ”Topics in classical automorphic forms,” G.S.M. AMS, Vol. 17 (1997).

FAMILIES AND SYMMETRY:

11. Katz and Sarnack, ”Zeros of zeta functions and symmetry,” BAMS, 36, 1-26 (1999).

12. Keating and Snaith, ”Random matrix theory and L-functions at S = \frac{1}{2},” Comm. Math Phys. 214, 91-110 (2000).

13. P. Michel, “Repartition des zeros des functions L et matrices aleatoires,” Seminar Bourbaki, March 2001.